OpenStudy (anonymous):

Which fraction is equivalent to 68%? A. B. C. D.

Nnesha (nnesha):

none :D

OpenStudy (demonchild99):

all you have to do is divide numerator by denomenator...and round to the nearest..u should get it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

A = 17/50 B = 17/25 C = 34/25 D = 680/100

OpenStudy (anonymous):

DIDNT PUT THE ANSWERS LOLOLOLOl

OpenStudy (anonymous):

@demonchild99 @Nnesha

OpenStudy (anonymous):

68/100 but if you want it simplified it would be 17/25

OpenStudy (anonymous):

B

OpenStudy (demonchild99):

yea

OpenStudy (alexandervonhumboldt2):

68%=68/100

OpenStudy (anonymous):

IM UR ALL FANS

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wait a medal?

OpenStudy (alexandervonhumboldt2):

i gave you

OpenStudy (anonymous):

how do i get a medal?

OpenStudy (anonymous):

or how do i give it?

OpenStudy (demonchild99):

i think it would be d

OpenStudy (alexandervonhumboldt2):

best responce

OpenStudy (anonymous):

d or b?

OpenStudy (demonchild99):

becuz it doesnt say to simplify

OpenStudy (anonymous):

press best response

OpenStudy (alexandervonhumboldt2):

b

OpenStudy (anonymous):

2 b's 1 d's

OpenStudy (demonchild99):

oop sorry my bad...b

OpenStudy (anonymous):

it wouldn't be 680 over 100

OpenStudy (anonymous):

of course not

OpenStudy (demonchild99):

i just realized that..lol

OpenStudy (anonymous):

wait

OpenStudy (demonchild99):

sorry

OpenStudy (anonymous):

@ICECREAMyummy

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

its ok

OpenStudy (triciaal):

percent means per 100 68% means 68/100 then find common factor to cancel out to reduce to lowest term

OpenStudy (demonchild99):

so i was right?

OpenStudy (triciaal):